OpenStudy (anonymous):

Can someone show the steps to solve this other equation? 2(3t+1)-5=t-(t+2)

OpenStudy (anonymous):

just multiply then solve for t 2(3t+1) will give 6t+2 , and t-(t+2) will give t-t-2 = -2 equalize them 6t+2=-2 6t=-4 t=-4/6

OpenStudy (anonymous):

I understood the 2(3t+1) part but i dont get the rest.

OpenStudy (anonymous):

try to read again, t-(t+2) will give t-t-2 = -2 now substitute them in the original formula 6t+2 (from 2(3t+1)) = -2 (from t-(t+2)) 6t+2=-2 6t=-4 t=-4/6

OpenStudy (anonymous):

6t+2-5=t-t-2 6t-3=-2 6t=3-2 6t=1 t=1/6

OpenStudy (anonymous):

man I forgot the minus 5 ><.. 6t+2-5=-2 6t-3=-2 6t=-2+3 6t=1 t=1/6

OpenStudy (anonymous):

6t+2=8t then -5 makes it 3t? How'd you get -2?

OpenStudy (anonymous):

i got -2 by distributing the negative there to (t+2) which gave me -t-2

OpenStudy (anonymous):

pat

OpenStudy (anonymous):

yes?